The length of a rectangle is 4 cm less that three times its width. If the length is increased by 8 cm and the width is decreased by 2 cm, the perimeter will be 100 cm. Find the dimensions of the orifinal rectangle.
Answer by stanbon(75887)

The length of a rectangle is 4 cm less that three times its width. If the length is increased by 8 cm and the width is decreased by 2 cm, the perimeter will be 100 cm. Find the dimensions of the orifinal rectangle.
---------------
Old Dimensions:
width=x
length = 3x-4
------------------
New Dimensions:
width = x-2
length = 3x-4+8=3x+4
---------
EQUATIOn:
Using new dimensions 2(length + width) = 100 cm
2(x-2+3x+4)=100
4x+2=50
4x=48
x=12
---------
Old width was 12 cm
Old length was 3(12)-4=32 cm
Cheers,
Stan H.